As a portfolio manager, you are deciding whether to add Crico stock to your portfolio. Crico Corporation is a publicly listed real estate investment trust. The current trading price of Crico’s stock is $175 per share. You have assembled the following financial information:

  1. The market value of its debt is $70 billion.
  2. The equity beta is .80; the market return is 4 percent; the risk-free rate is 2 percent.
  3. The company has 330 million shares outstanding
  4. The before-tax cost of debt is 6 percent.
  5. The tax rate is 35 percent.
  6. The company’s capital structure is 60 percent debt, and 40 percent common stock. There is no preferred stock.
  7. BCC’s FCFF for the year just ended is $1 billion. FCFF is expected to grow at a constant rate of 3 percent forever.

Expected cost of equity (using CAPM) = risk-free rate + beta*(market return - risk-free rate)

= 2% + 0.80*(4% - 2%) = 3.60%

What is the estimated weighted average cost of capital?
Answer #1

Capital Structure: Debt = 60 % and Common Stock = 40 %, Before-Tax Cost of Debt = 6% and Tax Rate = 35 %

After-Tax Cost of Debt = (1-Tax Rate) x Before-Tax Cost of Debt = kd = (1-0.35) x 6 = 3.9 %

Risk_free Rate = 2 % , Market Return = 4 % and Beta = 0.8

Cost of Equity = ke = 2 + 0.8 x (4-2) = 3.6 %

Weighted Average Cost of Capital (WACC) = kd x Debt Proportion + ke x Common Stock Proportion = 3.9 x 0.6 + 3.6 x 0.4 = 3.78 %